Find an interval where the following functions has a root. Show all you work how you chose the interval. i) f(x) = 3x + sin(x) βˆ’ e^x


1
Expert's answer
2021-08-19T07:03:56-0400

Use the Intermediate Value Theorem

The function "f(x)=3x+\sin(x)-e^x" is continuous on "(-\infin, \infin)." Then the function "f" is continuous on "[0,1]."

"f(0)=3(0)+\sin(0)-e^0=-1<0"

"f(1)=3(1)+\sin(1)-e^1>0"

By the Intermediate Value Theorem "\exist c\in(0,1)" such that "f(c)=0."

The function "f(x)=3x+\sin(x)-e^x" is continuous on "[1, 2]."


"f(2)=3(2)+\sin(2)-e^2<7-e^2<0"

By the Intermediate Value Theorem "\exist c\in(1,2)" such that "f(c)=0."



"f'(x)=3+\cos(x)-e^x"




"f'(x)=3+\cos(x)-e^x>3+\cos(x)-1>0, x\leq0"

"f'(x)=3+\cos(x)-e^x>3+\cos(x)-e>0, 0\leq x\leq1"

Hence "f(x)" strictly increases on "(-\infin, 1)"

Therefore the function "f" has the only root on "(-\infin, 1]."

Therefore we choose the interval "[0, 1]" to find the root of the function "f."



"f'(x)=3+\cos(x)-e^x\leq4-e^2<0, x\geq2"

Hence "f(x)" strictly decreases on "(2, \infin)"

Therefore the function "f" has no root on "[2, \infin)."

Therefore we choose the interval "[1, 2]" to find the root of the function "f."


Therefore we choose the interval "[0, 2]" to find all roots of the function "f."